# PSZ: Enhancing the SZ Scientific Lossy Compressor With Progressive Data Retrieval

### v1

@@ 6.2. Evaluation Results and Analysis
The evaluation covers three aspects – compression and decompression/retrieval speed, compression ratio, and progressive retrieval effectiveness under error-bound or bitrate constraints. To be more specific, the speed test confirms IPComp’s high compression efficiency, especially compared to residual-based alternatives (SZ3-R and ZFP-R) whose speed drops significantly when increasing the number of residuals. The compression ratio comparison demonstrates our solution IPComp leads to the smallest compressed data size. The progressive retrieval evaluation shows that IPComp requires the lowest data volume to reconstruct to the same fidelity compared with others, and IPComp leads to the highest reconstruction fidelity under the same bit rate budget.
6.2.1. Speed
We evaluate the compression and retrieval speed of all the compressors. To ensure a fair comparison, we use $eb=10^{-9}\times\text{Range}(dataset)$ to compress and retrieve the data for all compressors, except for PMGARD as it is a lossless compress with lossy retrieval by design. To reach the target $eb$, residual-based compressors need to compress and decompress in multiple iterations. The speed tests are shown in Figure 5. It confirms our solution IPComp is up to around 300% faster than other approaches in most cases. We include one additional compressor SPERR-R in this figure. SPERR-R is the residual progressive implementation of the wavelet-based lossy compressor SPERR (Li et al., 2023). The speed of SPERR-R is extremely slow – less than 50 MB/s in more than half of the cases. Such a slow speed will offset the savings of reduced data size for I/O, and potentially slow down the scientific workflow. As a result, we do not include SPERR-R as one of our baselines in the full evaluation.
(a) Compression
(b) Decompression
Figure 5. Our solution IPComp has the fastest speed in both compression and decompression compared with others
(a) Compression
(b) Decompression
Figure 6. The speed of residual-based progressive solutions (SZ3-R, ZFP-R) will drop significantly when the number of residuals increases
Another evaluation in this section is the speed of residual-based compressors. As we discussed in Section 6.1.3, residual-based ones need to compress and decompress multiple times based on the number of pre-defined error bounds. More pre-defined error bounds would provide more flexibility in retrieval, however, as Figure 6 shows, this will instead reduce their speed significantly. We also observe from Figure 6 that the residual results are curved instead of straight lines. The reason is that although having more predefined error bounds increases the number of iterations, each iteration takes less time because the looser bounds result in a smaller range of quantized integers. However, the total time still increases significantly due to the cumulative effect of all iterations. The drawback of residual-based compressors in speed highlights the advantages of our solution which can deliver highly flexible retrieval in high speed.
6.2.2. Compression Ratios
Compression ratio (cr) is a key metric for evaluating any compressor, regardless of whether it is progressive or not. As discussed in Section 1, the fundamental goal of compression is to reduce data size, as smaller data sizes alleviate the burden on applications in terms of storage, transfer, and processing.
(a) High precision setting ($eb=1e-9$)
(b) High ratio setting ($eb=1e-6$)
@@ 6.2.2. Compression Ratios
Figure 7 presents the compression ratios of all solutions under the same error-bound settings. We select two error bounds which are 1e-6 and 1e-9 to cover both high precision (cr¡10) and high ratio (cr up to 50) cases. As this figure shows, our solution IPComp achieves a compression ratio advantage of around 20% to 500% over other progressive compressors across the vast majority of datasets. This allows users to perform progressive decompression while utilizing the least possible storage space for compressed data.
In fact, although not shown in the figure, IPComp achieves an even higher compression ratio than the non-progressive SZ3 particularly in high-precision scenarios, despite both being based on the interpolation prediction algorithm. The primary reason for this improvement lies in the encoding process. The non-progressive SZ3 uses Huffman entropy encoding, which is then further compressed using Zstd. Since Huffman coding assigns variable-length codes based on frequency, it may disrupt certain repetitive patterns in the byte or word level after encoding. This, in turn, reduces the effectiveness of Zstd’s lossless compression, as Zstd relies on detecting and exploiting repetitive patterns at the byte/word level to achieve higher compression efficiency. In contrast, our solution IPComp employs a customized predictive encoding method, discussed in Section 4.4, followed by Zstd for entropy encoding and pattern extraction. By avoiding the disruption caused by Huffman and preserving more repetitive structures, our method enhances the final compression ratio.
@@ 6.2.3. Progressive retrieval efficiency
Figure 8 demonstrate the evaluation results of the error bound-based mode, where users specify a target error bound (max error tolerance). Compressors with higher efficiency should result in smaller data retrieval size under the error restrictions. The retrieval size is shown as bitrate in Figure 8, so lower lines in the figure indicate better results.
Figure 9 presents the results in fixed bitrate mode, where users set a target retrieval size. Compressors with higher efficiency should result in higher fidelity (indicated by lower error) with the same retrieval size, so lower lines in the figure indicate better results.
@@
Our solution is the only progressive compressor among the baselines that directly supports the fixed bitrate mode. As a result, manual efforts are needed to test those baselines. We enable fix bitrate mode for residual-based compressors by selecting the largest anchor point that fits within the user-specified bitrate constraint. Similarly, for PMGARD, we manually define anchor points ranging from $2^{16}eb$, $2^{15}eb$, $2^{14}eb$ down to $eb$, to allow for bitrate-based decompression.
@@
As shown in both Figure 8 and Figure 9, our compressor IPComp demonstrates its superior compression efficiency by consistently achieving the smallest data load size under the same max error, and the lowest max error under the same bitrate budget. This advantage primarily comes from the predictive coding method discussed in Section 4.4 and the optimized loading strategy discussed in Section 5. The residual-based solutions SZ3-R and ZFP-R have a staircase line in both of the figures because the fidelity of such a solution is limited to a few pre-defined residual levels. In comparison, our solution provides higher flexibility as it supports retrieval on arbitrary fidelity and bitrate.
@@ 6.2.4. Progressive retrieval for PSNR
We also evaluate the PSNR of the compressed data. Since PSNR is mathematically related to the $L_{2}$ norm, it serves as another important metric for assessing compression fidelity. Although our design primarily targets the $L_{\infty}$ norm and we do not explicitly optimize for PSNR, Figure 10 shows our approach still maintains competitive or superior PSNR compared to other baselines.
@@ 6.2.5. Visual quality
Additionally, we evaluate the visualization quality of the reconstructed data as shown in Figure 11. We load 0.1%, 0.3%, and 1.0% of the data and assess the impact on visualization of Curl and Laplacian metrics. While loading 0.3% of data is enough for Curl in terms of visualization, 1% of data is needed for Laplacian. This confirms the necessity of progressive retrieval in scientific applications.

### v2

@@ 6.2. Evaluation Results and Analysis
The evaluation covers three aspects – compression ratio, progressive retrieval effectiveness under error-bound or bitrate constraints, and compression and decompression/retrieval speed. To be more specific, the compression ratio comparison demonstrates our solution IPComp leads to the smallest compressed data size. The progressive retrieval evaluation shows that IPComp requires the lowest data volume to reconstruct to the same fidelity compared with others, and IPComp leads to the highest reconstruction fidelity under the same bit rate budget. The speed test confirms IPComp’s high compression efficiency, especially compared to residual-based alternatives (SZ3-R and ZFP-R) whose speed drops significantly when the number of residual compressions increases.
(a) High precision setting ($eb=1\mathrm{e}{-9}$)
(b) High ratio setting ($eb=1\mathrm{e}{-6}$)
@@
6.2.1. Compression Ratios
CR is a key metric for evaluating any compressor, regardless of whether it is progressive or not. As discussed in Section 1, the fundamental goal of compression is to reduce data size, as smaller data sizes alleviate the burden of storage, transfer, and processing on applications.
Figure 5 presents the compression ratios of all solutions under the same error-bound settings. We select two error bounds which are 1e-6 and 1e-9 to cover both high precision (cr¡10) and high ratio (cr up to 50) cases. As this figure shows, our solution IPComp achieves a compression ratio advantage of around 20% to 500% over other progressive compressors across the vast majority of datasets. This allows users to perform progressive decompression while utilizing the least possible storage space.
In fact, although not shown in the figure, IPComp achieves an even higher compression ratio than the non-progressive SZ3 particularly in high-precision scenarios, despite both being based on the interpolation prediction algorithm. The primary reason for this improvement lies in the encoding process. The non-progressive SZ3 uses Huffman entropy encoding, which is then further compressed using zstd. Since Huffman coding assigns variable-length codes based on frequency, it may disrupt certain repetitive patterns in the byte or word level after encoding. This, in turn, reduces the effectiveness of zstd’s lossless compression, as zstd relies on detecting and exploiting repetitive patterns at the byte/word level to achieve higher compression efficiency.  (Collet and Kucherawy, 2018) In contrast, our solution IPComp employs a customized predictive encoding method, discussed in Section 4.4, followed by zstd for entropy encoding and pattern extraction. By avoiding the disruption caused by Huffman and preserving more repetitive structures, our method enhances the final compression ratio.
@@ 6.2.2. Progressive retrieval efficiency
Figure 6 demonstrate the evaluation results of the error bound mode. Compressors with higher efficiency should result in smaller data retrieval size under the error restrictions. The retrieval size is shown as bitrate in Figure 6, so lower lines in the figure indicate better results.
@@
Figure 7 presents the results in fixed bitrate mode. Compressors with higher efficiency should result in higher fidelity (indicated by lower error) with the same retrieval size, still lower lines in the figure indicate better results. Since our solution is the only progressive compressor that directly supports the fixed bitrate mode, manual efforts are needed to test other baselines. We enable fix bitrate mode for residual-based compressors by selecting the largest anchor point that fits within the user-specified bitrate constraint. Similarly, for PMGARD, we manually define anchor points ranging from $2^{16}eb$, $2^{15}eb$, $2^{14}eb$ down to $eb$, to allow for bitrate-based decompression.
@@
As shown in both Figure 6 and Figure 7, our compressor IPComp demonstrates its superior compression efficiency by consistently achieving the smallest data load size under the same max error, and the lowest max error under the same bitrate budget. This advantage primarily comes from the predictive coding method discussed in Section 4.4 and the optimized loading strategy discussed in Section 5. The residual-based solutions SZ3-R and ZFP-R have a staircase line in both of the figures because the fidelity of such solutions is limited to a few pre-defined residual levels. In comparison, our solution provides higher flexibility as it supports retrieval on arbitrary fidelity and bitrate.
(a) Compression
(b) Decompression
Figure 8. Our solution IPComp has the fastest speed in both compression and decompression than all others except for SZ3-M which is multi fidelity but not progressive
(a) Compression
(b) Decompression
Figure 9. The speed of residual-based progressive solutions (SZ3-R, ZFP-R, etc) will drop significantly when the setting of residual count increases
6.2.3. Speed
We evaluate the compression and retrieval speed of all the compressors. To ensure a fair comparison, we use $eb=10^{-9}\times\text{Range}(dataset)$ to compress and retrieve the data for all compressors, except for PMGARD as it is a lossless compress with lossy retrieval by design. To reach the target $eb$, residual-based compressors need to compress and decompress in multiple iterations. The speed tests are shown in Figure 8. It confirms our solution IPComp is up to around 300% faster than other approaches in most cases, except for SZ3-M. SZ3-M supports multi fidelity by compressing input in different error bounds and storing all those outputs together. Such a solution is not progressive since the retrieval cannot use previous passes of data. As a result, its compression ratio and retrieval efficiency are both extremely limited in the experiments. We include one additional compressor SPERR-R in this figure. SPERR-R is the residual progressive implementation of the wavelet-based lossy compressor SPERR (Li et al., 2023). The speed of SPERR-R is extremely slow – less than 50 MB/s in more than half of the cases. Such a slow speed will offset the savings of reduced data size for I/O, and potentially slow down the scientific workflow. As a result, we do not include SPERR-R as one of our baselines in the full evaluation.
Another evaluation in this section is the speed of residual-based compressors. As discussed in Section 6.1.3, residual-based ones must compress and decompress multiple times based on the number of pre-defined error bounds. More pre-defined error bounds would provide more flexibility in retrieval, however, as Figure 9 shows, this will instead reduce their speed significantly. We also observe from Figure 9 that the residual results are curved instead of straight lines. The reason is that although having more predefined error bounds increases the number of iterations, each iteration takes less time because the looser bounds result in a smaller range of quantized integers. However, the total time still increases significantly due to the cumulative effect of all iterations. This drawback of residual-based compressors highlights the advantages of our solution which can deliver highly flexible retrieval at high speed.
@@ 6.2.4. Progressive retrieval for PSNR
We also evaluate the PSNR of the compressed data. Since PSNR is mathematically related to the $L_{2}$ norm error, it serves as another important metric for assessing compression fidelity. Although our design primarily targets the $L_{\infty}$ norm and we do not explicitly optimize for PSNR, Figure 10 shows our approach still maintains competitive or superior PSNR compared to other baselines.
@@ 6.2.5. Visual quality
Additionally, we evaluate the visualization quality of the reconstructed data as shown in Figure 11. We load 0.1%, 0.3%, and 1.0% from the same data, and assess the impact on visualization of Curl and Laplacian. While loading 0.3% of data is enough for Curl in terms of visualization, 1% of data is needed for Laplacian. This confirms the necessity of progressive retrieval in scientific applications.